# Model Tampering Attacks Enable More Rigorous Evaluations of LLM Capabilities

### v1

@@ Abstract
Evaluations of large language model (LLM) risks and capabilities are increasingly being incorporated into AI risk management and governance frameworks. Currently, most risk evaluations are conducted by designing inputs that elicit harmful behaviors from the system. However, a fundamental limitation of this approach is that the harmfulness of the behaviors identified during any particular evaluation can only lower bound the model’s worst-possible-case behavior. As a complementary method for eliciting harmful behaviors, we propose evaluating LLMs with model tampering attacks which allow for modifications to latent activations or weights. We pit state-of-the-art techniques for removing harmful LLM capabilities against a suite of 5 input-space and 6 model tampering attacks. In addition to benchmarking these methods against each other, we show that (1) model resilience to capability elicitation attacks lies on a low-dimensional robustness subspace; (2) the attack success rate of model tampering attacks can empirically predict and offer conservative estimates for the success of held-out input-space attacks; and (3) state-of-the-art unlearning methods can easily be undone within 16 steps of fine-tuning. Together these results highlight the difficulty of removing harmful LLM capabilities and show that model tampering attacks enable substantially more rigorous evaluations than input-space attacks alone. [^1] We release models at https://huggingface.co/LLM-GAT.
Machine Learning, ICML
Figure 1: Model tampering attacks modify latents and weights. In contrast to input-space attacks, model tampering attacks elicit capabilities from an LLM by making modifications to the internal activations or weights.
Figure 2: Pitting capability suppression (unlearning) methods against capability elicitation attacks. We evaluate unlearning methods and capability elicitation attacks on removing and extracting hazardous bio-knowledge. All unlearning methods tested could be successfully attacked. Left: The unlearning score (Equation 1) measures how effectively each unlearning method removed unwanted capabilities while preserving general model utility. Higher scores indicate better unlearning (scale 0-1). Right: Increase in the unlearned task performance after attacks. The first 5 columns are from input-space attacks while the final 6 are from model tampering attacks. In particular, finetuning attacks (rightmost columns) were especially effective at revealing suppressed capabilities.
@@ 1 Introduction: Limitations of Input-Output Evaluations
Rigorous evaluations of large language models (LLMs) are widely recognized as key for risk mitigation (Raji et al., 2022; Anderljung et al., 2023; Schuett et al., 2023; Shevlane et al., 2023) and are being incorporated into AI governance frameworks (NIST, 2023; UK DSIT, 2023; EU, 2023; China, 2023; Brazil, 2023; Canada, 2022; Korea, 2025). However, despite their efforts, developers often fail to identify overtly harmful LLM behaviors pre-deployment (Shayegani et al., 2023; Andriushchenko et al., 2024; Carlini et al., 2024; Yi et al., 2024b). Current methods primarily rely on automated input-space attacks, where evaluators search for prompts that elicit harmful behaviors. These are useful but often leave unidentified vulnerabilities. A difficulty with input-space attacks is that they are poorly equipped to cover the attack surface. This happens for two reasons. First, it is intractable to exhaustively search the input space. [^2] For example, with modern tokenizers, there are vastly more 20-token strings than particles in the known universe. Second, attackers can sometimes manipulate more than just model inputs (e.g., if a model is open-source). This highlights a fundamental limitation of input-space evaluations: the worst behaviors identified during an assessment can only offer a lower bound of the model’s overall worst-case behavior (Gal, 2024; OpenAI, 2024).
To address this challenge, we draw inspiration from a safety engineering principle: that safety-critical systems should be tested under stresses at least as extreme—if not more—than those expected in deployment (Clausen et al., 2006). For example, buildings are designed to withstand loads multiple times greater than their intended use. Here, we take an analogous approach to evaluating LLMs: stress-testing them under attacks that go beyond input-space manipulations.
@@
  • Question 2: Can model tampering attacks inform evaluators about LLM vulnerabilities to unforeseen input-space attacks? Answering this will help us understand how model tampering attacks can help assess risks from both open- and closed-source models.
To answer these questions, we pit state-of-the-art methods for unlearning and safety fine-tuning in LLMs against a suite of input-space and model tampering attacks (Figure 2). We make four contributions:
@@ 2 Related Work
Capability elicitation and evaluation: Research on adversarial capability elicitation (Hofstätter et al., 2024) in LLMs has primarily been done in the context of machine unlearning (Liu et al., 2024a; Barez et al., 2025) and jailbreaking (Yi et al., 2024b). Here, we experiment in these two domains. However, capability elicitation has also been researched in the context of backdoors/trojans (Zhao et al., 2024), “password-locked models” (Greenblatt et al., 2024), and “sandbagging” (van der Weij et al., 2024). In the unlearning field, several recent works have used adversarial methods to evaluate the robustness of unlearning algorithms (Patil et al., 2023; Lynch et al., 2024; Łucki et al., 2024; Hu et al., 2024; Liu et al., 2024a; Zhang et al., 2024; Liu et al., 2024b). Here, we build off of Li et al. (2024b) who introduce WMDP-Bio, a benchmark for unlearning dual-use biotechnology knowledge from LLMs. In the jailbreaking field, many techniques have been developed to make LLMs comply with harmful requests (Shayegani et al., 2023; Yi et al., 2024b; Jin et al., 2024; Chowdhury et al., 2024; Lin et al., 2024). Here, we experiment with 9 open-source LLMs and a set of gradient-guided, perplexity-guided, and prosaic techniques from the adversarial attack literature (see Table 1).
@@ 3 Methods
Weight space Benign LoRA Qi et al. (2023)
@@
Table 1: Table of capability elicitation (attack) and capability suppression (defense) defense. We consider defenses in two different settings: unlearning approaches that remove hazardous bio-knowledge and refusal-tuned models that resist jailbreaking attempts.
@@
Our approach: pitting capability suppression defenses against capability elicitation attacks. For unlearning experiments, we experiment with 65 models trained using 8 different unlearning methods. For jailbreaking experiments, we experiment with 9 models off the shelf from prior works. In both cases, we pit these defenses against a set of 11 input-space and model tampering attacks to either elicit ‘unlearned’ knowledge or jailbreak the model. In Table 1, we list all unlearning methods, off-the-shelf models, and attacks we use. Since the input-space attacks that we use are held out, we treat them as proxies for ‘unforeseen’ input-space attacks in our evaluations.
@@
Unlearned models for unlearning experiments: We unlearn dual-use bio-hazardous knowledge on Llama-3-8B-Instruct (Dubey et al., 2024) with the unlearning methods listed in Table 1 and outlined in Section A.2.1. For all methods, we train on 1,600 examples of max length 512 from the bio-remove-split of the WMDP ‘forget set’ (Li et al., 2024b), and up to 1,600 examples of max length 512 from Wikitext as the ‘retain set’. For the 8 unlearning methods listed in Table 1, we take 8 checkpoints evenly spaced across training. Finally, we also use the public release of the “TAR-v2” model from Tamirisa et al. (2024) as a 9th TAR model. In total, the 8 checkpoints each from the 8 methods we implemented plus the TAR model from Tamirisa et al. (2024) resulted in 65 models.
@@
Fine-tuned models for jailbreaking experiments: For jailbreaking experiments, we use the 9 fine-tuned Llama3-8B-Instruct models off the shelf listed in Table 1. The first 8 are all fine-tuned for robust refusal of harmful requests. Of these, ‘RR’ (Zou et al., 2023a) and ‘LAT’ (Sheshadri et al., 2024) are state-of-the-art for open-weight jailbreak robust models (Li et al., 2024a; Haize Labs, 2023). The final ‘Orenguteng’ model was fine-tuned to be ‘helpful-only’ and comply even with harmful requests. We discuss these models in more detail in Section A.3.
@@
Capability elicitation attacks: We use 5 input-space attacks and 6 model tampering attacks on our unlearned models. We use these attacks (single-turn) to increase dual-use bio knowledge (as measured by WMDP-Bio performance (Li et al., 2024b)) for unlearning experiments and to elicit compliance with harmful requests (as measured by the StrongReject AutoGrader (Souly et al., 2024)) for jailbreaking experiments. We selected attacks based on algorithmic diversity and prominence in the state of the art. We list all 11 attacks in Table 1. In all experiments, we produce universal adversarial attacks optimized to work for any prompt. This allows us to attribute improvements in performance to high-level capability elicitation rather than answer-forcing from the model (e.g., Fort (2023)). For descriptions and implementation details for each attack method, see Appendix A. Finally, we also used two proprietary attacks – one for unlearning experiments and one for jailbreaking experiments which we will describe in Section 4.
Data for attacks: All attacks, except fine-tuning attacks, used 64 examples for attack generation. For unlearning attacks, we used 64 held-out examples from the WMDP-Bio test set ($\sim$5of the evaluation set). For jailbreaking attacks, we used 64 held-out examples of compliance with harmful requests from Sheshadri et al. (2024). For benign fine-tuning attacks, we used WikiText (Merity et al., 2016). For malicious fine-tuning unlearning attacks, we use either the WMDP ‘bio retain’ or ‘forget’ sets and report averages across both. For malicious fine-tuning jailbreaking attacks, we use examples of compliance with harmful requests from Sheshadri et al. (2024). For details on attack configurations, including the number of examples, batch size, number of steps, and other hyper-parameters, see Table 4.
@@
Model tampering attacks are efficient. In Table 3, we show the number of forward and backward passes used in our implementations of attacks. Model tampering attacks were more efficient than state-of-the-art input-space attacks.
@@ 4 Experiments
Table 2: Benchmarking LLM unlearning methods: We report results for the checkpoint from each method with the highest unlearning score. We report original WMDP-Bio performance, worst-case WMDP-Bio performance after attack, and three measures of general utility: MMLU, MT-Bench, and AGIEval. For all benchmarks, the random-guess baseline is 0.25 except for MT-Bench/10 which is 0.1. Representation rerouting (RR) has the best unlearning score. No model has a WMDP-Bio performance less than 0.36 after the most effective attack. We note that Grad Diff and TAR models performed very poorly, often struggling with basic fluency.
As discussed in Section 1, we have two motivations, each corresponding to a different threat model. First, we want to directly evaluate robustness to model tampering attacks to better understand the risks of open-source, leaked, or API fine-tuneable LLMs. Second, we want to understand what model tampering attacks can tell us about unforeseen input-space attacks in order to study risks from all types of LLMs. Unfortunately, unforeseen attacks are, by definition, ones that we do not have access to. Instead, since the input-space attacks that we use are held out during fine-tuning, we treat them as proxies for ‘unforeseen’ input-space attacks.
4.1 Unlearning
@@ 4.1.1 Benchmarking Unlearning Methods
The ability to robustly remove harmful knowledge has been proposed as a key goal for designing safer AI (Barez et al., 2025). However, our results show that state-of-the-art unlearning methods fail to achieve this goal. Even the best-performing unlearning approaches leave models vulnerable to few-shot fine-tuning.
Unlearning evaluation: In our models, we evaluate unlearning efficacy on the WMDP-Bio QA evaluation task (Li et al., 2024b). We evaluate general utility using three benchmarks. First, we use MMLU (Hendrycks et al., 2020) and AGIEval (Zhong et al., 2023), which are based on asking LLMs multiple-choice questions. We then use MT-Bench (Bai et al., 2024) which is based on long answer questions and thus measures both knowledge and fluency. Because the goal of unlearning is to differentially decrease capabilities in a target domain, we calculate an “unlearning score” based on both unlearning efficacy and utility degradation. Given an original model $M$ and an unlearned model $M^{\prime}$, we calculate $S_{\text{unlearn}}(M^{\prime})$ with the formula:
@@
Figure 3: Three principal components explain 89% of the variation in attack success. Left: The proportion of explained variance for each principal component. Right: We display the first three principal components weighted by their eigenvalues. The first principal component shows a split between the two adversarial (LoRA, Full) fine-tuning attacks and all others.
Figure 4: Hierarchical clustering reveals natural groupings of attacks. Attacks tend to cluster by algorithmic type. However, benign model tampering attacks cluster with gradient-free input-space attacks.
@@ 4.1.2 Model robustness exists on a low-dimensional subspace
Weighted PCA: First, to understand the extent to which some attacks offer information about others, we analyze the geometry of attack successes across our 65 models. We perform weighted principal component analysis on the WMDP-Bio improvements achieved by all 11 attacks on all 65 model checkpoints. We first constructed a matrix $A$ with one row per model and one column per attack. Each $A_{ij}$ corresponds to the increase in WMDP-Bio performance in model $i$ under attack $j$. We then centered the data and multiplied each row $A_{i}$ by the square root of the unlearning score: $\sqrt{S_{\text{unlearn}}(A_{i})}$. This allowed for models to influence the analysis in proportion to their unlearning score.
@@
Method Upper-Bounded $\uparrow$ Attack Strength $\uparrow$
@@
Hierarchical clustering reveals distinct attack families. In Figure 4, we perform agglomerative clustering on attack success correlations. Algorithmically similar attacks tend to cluster together. However, adversarial finetuning attacks exhibit significant variation, even amongst each other. Finally we see that benign model tampering methods behave similarly to gradient-free input-space attacks.
4.1.3 Model tampering attacks empirically help to predict and upper bound the success of input-space attacks
@@
Fine-tuning attack successes empirically upper bound input-space attack successes. LoRA and Full fine-tuning performed differently on different attacks. However, together, the max of the two did as well or better than the best-performing input-space attack on 64 of 65 models. This suggests that model tampering attacks could be used to develop more cautious estimates of a model’s worst-case behaviors than other attacks.
@@
Figure 6: Model tampering attacks remain predictive for a proprietary attack from the UK AI Safety Institute. (a) In these experiments, correlations are weaker than with non- UK AISI attacks, but benign fine-tuning attacks continue to correlate with UK AISI input-space attack success. (b) Fine-tuning attacks still tend to exceed the success of input-space attacks, though less consistently than with the attacks from Figure 5.
@@
Figure 7: Fine-tuning efficiently undoes unlearning. We plot the heatmap of the best checkpoint for each method under fine-tuning attack. All fine-tuning experiments are done within 16 gradient steps, with 128 examples or fewer. All methods can be attacked to increase WMDP-Bio performance by 10 percentage points or more. All hyper-parameters are listed in Table 4.
@@
Model tamperings are predictive of the success of proprietary attacks from the UK AI Safety Institute (UK AISI). To more rigorously test what model tampering attacks can reveal about unforeseen input-space attacks, we analyze their predictiveness on proprietary attacks from the UK AI Safety Institute. These attacks were known to the ‘red team’ authors ( UK AISI affiliates) but were not accessible to all other ‘blue team’ authors. We conducted these attacks with the same data and approach as all of our other input-space attacks. Results are summarized in Figure 6 with full results in the Appendix. Correlations are weaker than before, but pruning and benign fine-tuning still correlate with attack success with a $p$ value near zero. Also as before, fine-tuning attack successes often tend to be as strong or stronger than input-space attacks. However, this trend was weaker, only occurring for 60 of the 65 models. See Section B.2 for full results and further analysis of UK AISI evaluations.
@@
Model tampering attacks improve worst-case vulnerability estimation. Finally, we test if model tampering attacks offer novel information that can be used to predict worst-case behaviors better than input-space attacks alone (Figure 13 in Section C.1). We train linear regression models to predict worst-case input-space attack success rates with information from either (1) only input-space attacks, or (2) both input-space and model tampering attacks. We find that including model tampering attacks improves predictiveness (e.g. from 0.77 to 0.82 with three predictors). The best-performing combinations typically include attacks from multiple families, suggesting diverse attacks provide complementary signals by probing different aspects of model robustness.
@@
State-of-the-art unlearning can reliably be reversed within 16 fine-tuning steps – sometimes in a single step. We show the results of multiple fine-tuning attacks against the best-performing model from each unlearning method in Figure 7. All finetuning experiments, as detailed in Table 4, are performed within 16 gradient steps and with 128 or fewer examples. The only method that was resistant to few-shot fine-tuning attacks was TAR, in which only 1 out of the 9 fine-tuning attacks were able to increase the WMDP-Bio performance by over 10 percentage points. However, TAR models had low unlearning scores due to poor general utility, which renders their robustness to fine-tuning unremarkable. All utility-preserving state-of-the-art unlearning methods can be attacked successfully to recover more than 30 percentage points of WMDP performance. Moreover, the ‘Full-4’ attack does only a single gradient step (with a batch size of 64) and still increases the WMDP performance on 6 of the 8 methods by over 25 percentage points.
4.2 Jailbreaking

### v2

@@ Abstract
Evaluations of large language model (LLM) risks and capabilities are increasingly being incorporated into AI risk management and governance frameworks. Currently, most risk evaluations are conducted by designing inputs that elicit harmful behaviors from the system. However, this approach suffers from two limitations. First, input-output evaluations cannot evaluate realistic risks from open-weight models. Second, the behaviors identified during any particular input-output evaluation can only lower-bound the model’s worst-possible-case input-output behavior. As a complementary method for eliciting harmful behaviors, we propose evaluating LLMs with model tampering attacks which allow for modifications to latent activations or weights. We pit state-of-the-art techniques for removing harmful LLM capabilities against a suite of 5 input-space and 6 model tampering attacks. In addition to benchmarking these methods against each other, we show that (1) model resilience to capability elicitation attacks lies on a low-dimensional robustness subspace; (2) the attack success rate of model tampering attacks can empirically predict and offer conservative estimates for the success of held-out input-space attacks; and (3) state-of-the-art unlearning methods can easily be undone within 16 steps of fine-tuning. Together these results highlight the difficulty of suppressing harmful LLM capabilities and show that model tampering attacks enable substantially more rigorous evaluations than input-space attacks alone. [^1] We release 64 models at https://huggingface.co/LLM-GAT.
@@ 1 Introduction: Limitations of Input-Output Evaluations
Rigorous evaluations of large language models (LLMs) are widely recognized as key for risk mitigation (Raji et al., 2022; Anderljung et al., 2023; Schuett et al., 2023; Shevlane et al., 2023) and are being incorporated into AI governance frameworks (NIST, 2023; UK DSIT, 2023; EU, 2023; China, 2023; Brazil, 2023; Canada, 2022; Korea, 2025). However, despite their efforts, developers often fail to identify overtly harmful LLM behaviors pre-deployment (Shayegani et al., 2023; Andriushchenko et al., 2024; Carlini et al., 2024; Yi et al., 2024b). Current methods primarily rely on automated input-space attacks, where evaluators search for prompts that elicit harmful behaviors. These are useful but often leave unidentified vulnerabilities. A difficulty with input-space attacks is that they are poorly equipped to cover the attack surface. This happens for two reasons. First, attackers can sometimes manipulate more than just model inputs (e.g., if a model is open-source). Second, it is intractable to exhaustively search the input space. [^2] For example, with modern tokenizers, there are vastly more 20-token strings than particles in the known universe. These challenges highlight a fundamental limitation of input-space evaluations: the worst behaviors identified during an assessment can only offer a lower bound of the model’s overall worst-case behavior (Gal, 2024; OpenAI, 2024).
To help address this challenge, we draw inspiration from a safety engineering principle: that safety-critical systems should be tested under stresses at least as extreme—if not more—than those expected in deployment (Clausen et al., 2006). For example, buildings are designed to withstand loads multiple times greater than their intended use. Here, we take an analogous approach to evaluating and building safety cases (Clymer et al., 2024) for LLMs: stress-testing them under attacks that go beyond input-space manipulations.
Figure 1: Model tampering attacks modify latents and weights. In contrast to input-space attacks, model tampering attacks elicit capabilities from an LLM by making modifications to the internal activations or weights. In this paper, we use model tampering attacks to (1) directly evaluate risks from malicious tampering with open-weight models and (2) indirectly evaluate difficult-to-foresee input-space vulnerabilities in models.
@@
Question 2: Can model tampering attacks inform evaluators about LLM vulnerabilities to novel input-space attacks? Answering this will help us understand how model tampering attacks can help assess risks from both open- and closed-source models.
To answer these questions, we pit state-of-the-art methods for unlearning and safety fine-tuning in LLMs against a suite of input-space and model tampering attacks. We make four contributions:
@@ 2 Related Work
Capability elicitation and evaluation: Research on adversarial capability elicitation (Hofstätter et al., 2025) in LLMs has primarily been done in the context of machine unlearning (Liu et al., 2024a; Barez et al., 2025) and jailbreaking (Yi et al., 2024b). Here, we experiment in these two domains. However, capability elicitation has also been researched in the context of backdoors/trojans (Zhao et al., 2024), “password-locked models” (Greenblatt et al., 2024; Hofstätter et al., 2025), and “sandbagging” (van der Weij et al., 2024). In the unlearning field, several recent works have used adversarial methods to evaluate the robustness of unlearning algorithms (Patil et al., 2023; Lynch et al., 2024; Łucki et al., 2024; Hu et al., 2024; Liu et al., 2024a; Zhang et al., 2024; Liu et al., 2024b). Here, we build off of Li et al. (2024b) who introduce WMDP-Bio, a benchmark for unlearning dual-use biotechnology knowledge from LLMs. In the jailbreaking field, many techniques have been developed to make LLMs comply with harmful requests (Shayegani et al., 2023; Yi et al., 2024b; Jin et al., 2024; Chowdhury et al., 2024; Lin et al., 2024). Here, we experiment with 9 open-source LLMs and a set of gradient-guided, perplexity-guided, and prosaic techniques from the adversarial attack literature (see Table 1).
@@ 3 Methods
Benign LoRA Qi et al. (2023)
@@
Table 1: Table of capability elicitation (attack) and capability suppression (defense) defense. We consider defenses in two different settings: (top) unlearning approaches that remove hazardous bio-knowledge and (bottom) refusal-tuned models that resist jailbreaks.
@@
Our approach: pitting capability suppression defenses against capability elicitation attacks. For unlearning experiments, we experiment with 65 models trained using 8 different unlearning methods. For jailbreaking experiments, we experiment with 9 models off the shelf from prior works. In both cases, we pit these defenses against a set of 11 input-space and model tampering attacks to either elicit ‘unlearned’ knowledge or jailbreak the model. In Table 1, we list all unlearning methods, off-the-shelf models, and attacks we use. Since the input-space attacks that we use are held out, we treat them as proxies for novel input-space attacks in our evaluations (see also Hofstätter et al. (2025)).
@@
Defenses: machine unlearning methods: We unlearn dual-use bio-hazardous knowledge on Llama-3-8B-Instruct Dubey et al. (2024) with the unlearning methods listed in Table 1 and outlined in Section A.2.1. For all methods, we train on 1,600 examples of max length 512 from the bio-remove-split of the WMDP ‘forget set’ (Li et al., 2024b), and up to 1,600 examples of max length 512 from Wikitext as the ‘retain set’. For the 8 unlearning methods listed in Table 1, we take 8 checkpoints evenly spaced across training. Finally, we also use the public release of the “TAR-v2” model from Tamirisa et al. (2024) as a 9th TAR model. In total, the 8 checkpoints each from the 8 methods we implemented plus the TAR model from Tamirisa et al. (2024) resulted in 65 models.
@@
Defenses: refusal fine-tuned models: For jailbreaking experiments, we use the 9 fine-tuned Llama3-8B-Instruct models off the shelf listed in Table 1. The first 8 are all fine-tuned for robust refusal of harmful requests. Of these, ‘RR’ (Zou et al., 2023a) and ‘LAT’ (Sheshadri et al., 2024) are state-of-the-art for open-weight jailbreak robust models (Li et al., 2024a; Haize Labs, 2023). The final ‘Orenguteng’ model was fine-tuned to be ‘helpful-only’ and comply even with harmful requests. We discuss these models in more detail in Section A.3.
@@
Attacks: capability elicitation methods: We use 5 input-space attacks and 6 model tampering attacks on our unlearned models. We use these attacks (single-turn) to increase dual-use bio knowledge (as measured by WMDP-Bio performance (Li et al., 2024b)) for unlearning experiments and to elicit compliance with harmful requests (as measured by the StrongReject AutoGrader (Souly et al., 2024)) for jailbreaking experiments. We selected attacks based on algorithmic diversity and prominence in the state of the art. We list all 11 attacks in Table 1. In all experiments, we produce universal adversarial attacks optimized to work for any prompt. This allows us to attribute attack success to capability elicitation rather than answer-forcing from the model (e.g., Fort (2023)). For descriptions and implementation details for each attack method, see Appendix A. Finally, we also used two proprietary attacks – one for unlearning experiments and one for jailbreaking experiments which we will describe in Section 4.
Attacks: data
  • Attacks on unlearning – non-fine-tuning: we used 64 held-out examples of multiple-choice biology questions from the WMDP-Bio test set.
  • Attacks on unlearning – adversarial fine-tuning: we use the WMDP ‘bio retain’ or ‘forget’ sets. Both of which are comprised of biology papers.
  • Attacks on refusal training – all except benign fine-tuning: we used held-out examples of compliance with harmful requests from Sheshadri et al. (2024). Each example is a prompt + response pair.
  • Attacks on unlearning and refusal training – benign fine-tuning: For all benign fine-tuning attacks, we used WikiText Merity et al. (2016).
For details on attack configurations, including the number of examples, batch size, number of steps, and other hyper-parameters, see Table 4.
@@
Attacks: model tampering attacks are efficient. In Table 3, we show the number of forward and backward passes used in our implementations of attacks. Model tampering attacks were more efficient than state-of-the-art input-space attacks.
@@ 4 Experiments
Table 2: Benchmarking LLM unlearning methods: We report results for the checkpoint from each method with the highest unlearning score (Equation 1). We report original WMDP-Bio performance, worst-case WMDP-Bio performance across our attacks, and three measures of general utility: MMLU, MT-Bench, and AGIEval. Representation rerouting (RR) has the best unlearning score. No model has a WMDP-Bio performance less than 0.36 after the most effective attack. We note that Grad Diff and TAR models performed very poorly, often struggling with basic fluency.
As discussed in Section 1, we have two motivations, each corresponding to a different threat model. First, we want to directly evaluate robustness to model tampering attacks to better understand the risks of open-source, leaked, or API fine-tuneable LLMs. Second, we want to understand what model tampering attacks can tell us about novel, unforeseen input-space attacks in order to study risks from all types of LLMs. Unfortunately, unforeseen attacks are, by definition, ones that we do not have access to. Instead, since the input-space attacks that we use are held out during fine-tuning, we treat them as proxies for ‘unforeseen’ input-space attacks.
4.1 Unlearning Experiments
@@ 4.1.1 Benchmarking Unlearning Methods
Figure 2: Pitting capability suppression (unlearning) methods against capability elicitation attacks. We use unlearning methods to suppress bio-hazardous knowledge from LLMs and pit these against capability elicitation attacks seeking to re-elicit the unlearned knowledge. All unlearning methods tested could be successfully attacked. Left: The unlearning score (Equation 1) measures how effectively each unlearning method removed unwanted capabilities while preserving general model utility. Higher scores indicate better unlearning (scale 0-1). Right: Increase in the unlearned task performance after attacks. The first 5 columns are from input-space attacks while the final 6 are from model tampering attacks. In particular, finetuning attacks (rightmost columns) were especially effective at resurfacing suppressed capabilities.
Calculating an unlearning score: In our models, we evaluate unlearning efficacy on the WMDP-Bio QA evaluation task (Li et al., 2024b). We evaluate general utility using three benchmarks. First, we use MMLU (Hendrycks et al., 2020) and AGIEval (Zhong et al., 2023), which are based on asking LLMs multiple-choice questions. We then use MT-Bench (Bai et al., 2024) which is based on long answer questions and thus measures both knowledge and fluency. Because the goal of unlearning is to differentially decrease capabilities in a target domain, we calculate an “unlearning score” based on both unlearning efficacy and utility degradation. Given an original model $M$ and an unlearned model $M^{\prime}$, we calculate $S_{\text{unlearn}}(M^{\prime})$ with the formula:
@@
Figure 3: Three principal components explain 89% of the variation in attack success. Left: The proportion of explained variance for each principal component. Right: We display the first three principal components weighted by their eigenvalues. The first principal component suggests a geometric distinction between the two adversarial (LoRA, Full) fine-tuning attacks and all others.
Figure 4: Hierarchical clustering reveals groupings of attacks. Attacks tend to cluster by algorithmic type. However, benign fine-tuning attacks cluster with gradient-free input-space attacks.
@@ 4.1.2 Model robustness exists on a low-dimensional subspace
We perform PCA, weighting models by their unlearning score. First, to understand the extent to which some attacks offer information about others, we analyze the geometry of attack successes across our 65 models. We perform weighted principal component analysis on the WMDP-Bio improvements achieved by all 11 attacks on all 65 model checkpoints. We first constructed a matrix $A$ with one row per model and one column per attack. Each $A_{ij}$ corresponds to the increase in WMDP-Bio performance in model $i$ under attack $j$. We then centered the data and multiplied each row $A_{i}$ by the square root of the unlearning score: $\sqrt{S_{\text{unlearn}}(A_{i})}$. This allowed for models to influence the analysis in proportion to their unlearning score.
@@
Method Beaten $\uparrow$ Attack Strength $\uparrow$
@@
Hierarchical clustering reveals distinct attack families. In Figure 4, we perform agglomerative clustering on attack success correlations. Algorithmically similar attacks tend to cluster together. However, adversarial finetuning attacks exhibit significant variation, even amongst each other. Finally we see that benign model tampering methods (pruning and benign fine-tuning) behave similarly to gradient-free input-space attacks.
4.1.3 Model tampering attacks empirically predict and conservatively estimate the success of input-space attacks
@@
Fine-tuning attack successes empirically offer conservative estimates for input-space attack successes. LoRA and Full fine-tuning performed differently on different attacks. However, together, the max of the two did as well or better than the best-performing input-space attack on 64 of 65 models. This suggests that model tampering attacks could be used to develop more cautious estimates of a model’s worst-case behaviors than other attacks.
@@
Figure 6: Model tampering attacks are predictive for a held-out proprietary attack from the UK AI Security Institute. Each point corresponds to a model. (a) In these experiments, correlations are weaker than with non- UK AISI attacks, but benign fine-tuning attacks continue to correlate with UK AISI input-space attack success. (b) Fine-tuning attacks still tend to exceed the success of input-space attacks, though less consistently than with the attacks from Figure 5.
@@
Figure 7: Few-shot fine-tuning efficiently undoes unlearning. We plot the heatmap of the best checkpoint for each method under benign (left), LoRA (middle), and full-parameter (right) fine-tuning attacks. All fine-tuning experiments are done within 16 gradient steps, with 128 examples or fewer. All methods can be attacked to increase WMDP-Bio performance by 10 percentage points or more. All hyper-parameters are listed in Table 4.
@@
Model tampering attacks are predictive of the success of proprietary attacks from the UK AI Security Institute (UK AISI). To more rigorously test what model tampering attacks can reveal about novel input-space attacks, we analyze their predictiveness on proprietary attacks from the UK AI Security Institute. These attacks were known to the ‘red team’ authors ( UK AISI affiliates) but were not accessible to all other ‘blue team’ authors. We conducted these attacks with the same data and approach as all of our other input-space attacks. Results are summarized in Figure 6 with full results in the Appendix. Correlations are weaker than before, but pruning and benign fine-tuning still correlate with attack success with a $p$ value near zero. Also as before, fine-tuning attack successes often tend to be as strong or stronger than input-space attacks. However, this trend was weaker, only occurring for 60 of the 65 models. See Section B.2 for full results and further analysis of UK AISI evaluations.
@@
Model tampering attacks improve worst-case input-space vulnerability estimation. Finally, we test if model tampering attacks offer novel information that can be used to predict worst-case behaviors better than input-space attacks alone (Figure 13 in Section C.1). We train linear regression models to predict worst-case input-space attack success rates with information from either (1) only input-space attacks, or (2) both input-space and model tampering attacks. We find that including model tampering attacks improves predictiveness (e.g. from $r=0.77$ to $0.83$ with four predictors). The best-performing combinations typically include attacks from multiple families, suggesting diverse attacks provide complementary signals by probing different aspects of model robustness.
@@
State-of-the-art unlearning can reliably be reversed within 16 fine-tuning steps – sometimes in a single step. We show the results of multiple fine-tuning attacks against the best-performing model from each unlearning method in Figure 7. All finetuning experiments, as detailed in Table 4, are performed within 16 gradient steps and with 128 or fewer examples. The only method that was resistant to few-shot fine-tuning attacks was TAR, in which only 1 out of the 9 fine-tuning attacks were able to increase the WMDP-Bio performance by over 10 percentage points. However, TAR models had low unlearning scores due to poor general utility, which renders their robustness to fine-tuning unremarkable. All utility-preserving state-of-the-art unlearning methods can be attacked successfully to recover more than 30 percentage points of WMDP performance. Moreover, even when we perform a single gradient step (with a batch size of 64) still increases the WMDP performance on 6 of the 8 methods by over 25 percentage points (see column “Full-4” in Figure 7).
4.2 Jailbreaking Experiments